Session: Oct 2011 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

JIVE (Jun Yang at Mon Oct 17 08:43:50 2011) Success:
This is a test message for the EVN feedback

Westerbork (Geert Kuper at Thu Oct 20 15:08:27 2011) Success:
Observing ok. FS logfile has bad values for outside temperature and humidity due to broken sensor.

Medicina (Roda at Thu Oct 20 15:11:09 2011) Success:

Medicina (Andrea Orlati at Thu Oct 20 15:35:05 2011) Minor failures:
VC07 probably swapped polarization with VC08

Urumqi (Zhang hua at Fri Oct 21 02:59:36 2011) Success:

Onsala (Michael Lindqvist at Fri Oct 21 06:19:19 2011) Failure: WIND
Could only observe for an hour. To windy after that.

Hartebeesthoek (Jonathan Quixk at Fri Oct 21 06:54:06 2011) Success:
One FTP transfer failed but was manually restarted. No other known problems

Shanghai (Bo Xia at Fri Oct 21 08:12:15 2011) Success:

Effelsberg (Uwe Bach at Fri Oct 21 11:06:42 2011) Success:
all well!

Yebes (Susana García-Espada at Fri Nov 04 09:23:59 2011) Success:
Not known problems.

Torun (Magda Kunert-Bajraszewska at Mon Nov 14 09:42:22 2011) Success: